how do i show my work for this problem step by step here it is:
e^ln5

Answer by stanbon!(97)

Let y=e^(ln5)
Take the ln of both sides to get:
lny = (ln5)(lne)
But lne=1, therefore
lny = ln5
Take the anti-log of both sides to get:
y=5
Therefore e^(ln5) = 5
Cheers,
Stan H.